Mathematics 134 – Intensive Calculus for Science 2 Quiz 7 – March 31, 2006

Your Name:

Do all work on this sheet (front and back). You have 15 minutes. No calculators on this quiz.

A) (10) For which value(s) of the constant c is y = c + e^{4x} a solution of \frac{dy}{dx} = 4y − 8?

B) (10) The following plot shows the slope field for the differential equation \frac{dy}{dx} = 3x − 2y. On the plot, roughly sketch in the graphs of the solutions with y(0) = 2 and y(0) = −2.

Please turn over for question C C) (5) Do two steps of Euler’s Method with Δx = .5 to approximate the solution of \frac{dy}{dx} = x + 2y with y(0) = 1.